OpenStudy (anonymous):

How do you find the numbers in the equation~ 4X+3Y=24? Thnx

OpenStudy (anonymous):

do u mean hw to find the value of x and y???

OpenStudy (anonymous):

yep, do u know how

OpenStudy (anonymous):

this equation has an infinite number of solutions in order to have only one solution for eah x and y u need two such equations

OpenStudy (anonymous):

1 sec got to find it

OpenStudy (anonymous):

it's 3X+4Y=24

OpenStudy (anonymous):

3y=24-4x y=8-4x/3

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

4x=24-3y x=6-3x/4

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

x will be equal to 6-3y/4

OpenStudy (anonymous):

Thank you guys so much!